OpenStudy (anonymous):

Solve the logarithmic equation. Be sure to reject any value that is not in the domain of the original logarithmic expressions. Give the exact answer.

OpenStudy (anonymous):

\[\ln \sqrt{x+6}=9\]

OpenStudy (anonymous):

ln (x+6)^(1/2) = 9 ln (x+6) = 18 Exponentiate -> x+6 = e^18 -> x = e^18-6

OpenStudy (anonymous):

Thanks!!!

OpenStudy (anonymous):

u r welcome.